OpenStudy (anonymous):

Sal opens a box of a dozen chocolate cremes and generously offers two of them to his friend Val. Val likes vanilla cremes the best, but all of the chocolates look alike on the outside. If four of the cremes are vanilla, what is the probability that both of Val’s picks turn out to be vanilla?

OpenStudy (anonymous):

For the first selection , the probability is \[\frac{ 4}{12}. \] After the first election there are 11 left . So the second probability becomes , \[ \frac{ 3 }{11} \] Now multiply each of the above probabilities to obtain : \[\frac{ 4 }{12 }*\frac{3 }{ 11 }= \frac{ 12 }{ 132 }\\ =\frac{ 1}{11 } \]

OpenStudy (anonymous):

kk